OpenStudy (anonymous):

Write the equation of the line that passes through (-1, 4) with a slope of -1 in standard form. A) -x + 4y = -1 B) x - 4y = -1 C) x + y = 3 D) x - y = -3

OpenStudy (anonymous):

first of all, c is the answer I'll write the step down

OpenStudy (anonymous):

Thank you! And yes, please show me how to do it. :)

OpenStudy (anonymous):

y=mx+b slope: -1 y=-x+b coordinate: (-1,4) 4=-1(-1)+b 4=1+b b=3 y=-x+3 y+x=3 x+y=3 (C)

OpenStudy (anonymous):

Can you explain what the mx + b stands for? Like where does slope go and all the others go?

OpenStudy (anonymous):

y=mx+b is a slope intercept form. m stands for slope and b stands for y-intercept

OpenStudy (anonymous):

Oh ok thanks!!! :)

OpenStudy (anonymous):

sure